OpenStudy (anonymous):

find the complex zeros of x^3+27. write f in factored form

OpenStudy (anonymous):

Difference of 2 cubes (x+3)(x^2-3x+9)=0 We know one zero is -3, you can apply the quadratic formula to x^2-3x+9 to find the 2 complex zeros

OpenStudy (anonymous):

\[x=\frac{3 \pm \sqrt{9-4(1)(9)}}{2(1)}\] \[x=\frac{3 \pm 3i \sqrt{3}}{2}\]

OpenStudy (anonymous):

\[f(x)=(x+3)(x-\frac{3-3i \sqrt{3}}{2})(x-\frac{3+3i \sqrt{3}}{2})\]

OpenStudy (anonymous):

i said difference of 2 cubes above, i meant sum of 2 cubes, but i applied the right formula.

OpenStudy (anonymous):

thank you